OpenStudy (anonymous):

Help please! The hypotenuse of a right triangle is 41 in long. One leg of the same triangle is 40 in long. What is the length of the other leg? Explain how you found your answer.

563blackghost (563blackghost):

well a^2+b^2=c^2 so we input those 40^2+b^2=41^2

OpenStudy (anonymous):

So b = 9?

563blackghost (563blackghost):

correct

OpenStudy (anonymous):

Alright, I got it. Thank you both for your help!!

563blackghost (563blackghost):

np xD